OpenStudy (anonymous):

You have three $1 bills, four $5 bills, and two $10 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill at random. Find P($5 then $10).

OpenStudy (anonymous):

Answer: you got more money than me xD

OpenStudy (mathstudent55):

For the first pick, how many bills are there in total? For the first pick, how many $5 are there? For the second pick, how many bills are there in total? For the second pick, how many $10 are there?

OpenStudy (anonymous):

Lol, Mttblink. and hmmm 1-9 2-4 3-9 4-2 i think?

OpenStudy (mathstudent55):

1 and 2 are correct. there are a total of 9 bills and there are 4 $5 bills before you pick the first time. For 3 and 4, once you already picked a $5 bill, and you are ready to do the second pick, there are now only 8 bills left, not 9. Of those, 2 are $10 bills. OK?

OpenStudy (anonymous):

okay. i need it in fraction form though

OpenStudy (mathstudent55):

The probability of the the 2 picks are: P($5) = 4/9 P($10) = 2/8 = 1/4 So P($5 then $10) = P($5) * P($10)

OpenStudy (anonymous):

I dont uderstand :c

OpenStudy (mathstudent55):

Multiply the two probabilities. P($5 then $10) = P($5) * P($10) = 4/9 * 1/4 = 1/9

OpenStudy (anonymous):

Is the answer 2 over 27?